Apply Cardano’s method for finding the roots of .0 2x 3x
1
Expert's answer
2019-04-09T10:02:35-0400

To apply the Cardano method it is necessary to bring this equation to an canonical form:

y^3+py+q=0


write down our cubic equation in the general form

ax^3+bx^2+cx+d=0

1*x^3+1*x^2+0*x+0=0


We will replace the variable


x=y-(b/3a)

by changing the variable can be reduced to the above canonical form with coefficients

p=c/a-(b^2/3a^2)=-1/3

q= (2b^3/27a^3)-(bc/3a^2)+d/a=2/27


the canonical form will have this form

y^3-1/3y+2/27=0


Determine the value

Q=(p/3)^3+(q/2)^2=-1/729+1/729=0

Q = 0 is a single real real root


y=a+b


where

a=((-q/2)+Q^½)^⅓

b=((-q/2)-Q^½)^⅓


y=(-1/3)+(-1/3)=-2/3


then

x= (-2/3)-(1/3)=-1